6/800 as a percent
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

6/800 as a percent
---------
To get percent, multiply by 100
6/800 = 600/800%
= 3/4 % or 0.75%

Answer by stanbon(75887)

6/800 as a percent
------------------
Divide numerator and denominator by 8 to get:
(6/8)/100
= (3/4)/100
= 0.75/100
= 0.75%
=============
Cheers,
Stan H.